1. Jo Ann and her lab partner are completing a free fall lab in physics.



Her partner drops the ball in front of the motion sensor to measure



its speed. Assuming acceleration is constant and air resistance is



negligible, how far does the object fall if it reaches a speed of 9.8



m/s? (3 points)



2. Anneth and her partner disagree about whether or not she can make



a picket fence drop to the ground in exactly 3.0 s. Her partner thinks



the fence is too light to drop that quickly, but she’s convinced it’s possible. Assuming acceleration is constant and air resistance is



negligible, from how high does she have to drop the picket fence for



the smart gate to measure its ground impact at exactly 1.0 s?



1
Expert's answer
2022-04-29T06:09:38-0400

1.

"s=\dfrac{gt^2}{2}, v=gt"

Then

"s=\dfrac{v^2}{2g}=\dfrac{(9.8m/s)^2}{2(9.8m/s^2)}=4.9m"

2.


"h=\dfrac{gt^2}{2}=\dfrac{9.8m/s^2(1s)^2}{2}=4.9m"